40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12995     Entry time: Wed May 17 08:19:59 2017
Author: Steve 
Type: Update 
Category: SUS 
Subject: 4.1M earthquake  

Sus dampings recovered. ETMY oplev needs to be recentered.

GV May 17 11am: I shut down the BS, SRM, ITMX and ITMY watchdogs, as the coil-driver boards for these optics are presently not installed.
 

Attachment 1: eq_4.1_SantaBarbara.png  829 kB  | Hide | Hide all
eq_4.1_SantaBarbara.png
Attachment 2: 4.1m_Isla_Vista_CA.png  56 kB  Uploaded Wed May 17 09:45:11 2017  | Hide | Hide all
4.1m_Isla_Vista_CA.png
ELOG V3.1.3-